# Airborne Sense and Detect of Drones using Deep Learning and LiDAR Point Clouds

### v2

@@ 4.2 Data Augmentation
4.2.1 LiDAR simulation with NVIDIA Optix Ray Tracing
@@
  • Step 2. Move the origin of the coordinate system to the centre of the drone mesh model.
@@
Each OptiX thread computes a new ray origin and direction vector independently, rather than rebuilding the acceleration structure in ray tracing core.
@@
4.3 Training Parameter
The parameters for the training process are as follows. For a single GPU, we limit the batch size to 4 due to memory constraints. However, when training with 8 GPUs, we increase the batch size to 32. During our experiments, we noticed that after 128 epochs, the accuracy did not improve further. Regarding the learning rate, we deviate from the original PointPillars paper (which used 0.0001) and set it to 0.001. Our empirical observations indicate that this value is more suitable for our use case. Safety is critical in our use case scenario, so we avoid any point cloud sampling. Otherwise, small objects like drones might be excluded from the scene. In our use case, we have not encountered any situations where sampling out occurs, when we set the maximum number of points each pillar to 100 and the maximum number of pillars to 12,000. However, these two values are changeable, they depend on the use case. To determine the optimal threshold values for positive IoU and negative IoU, we conducted numerous experiments with various combinations. Our observations indicate that the combination of 0.4 for positive IoU and 0.35 for negative IoU yields the highest accuracy.
5 Model Test
5.1 Model Test with Digital Twin
The LiDAR point data is accumulated into time windowed frames and then input to the detection model. The optimal time window depends on several factors. The upper threshold is determined by the acceptable computational latency which increases in proportion to the number of points in each frame, while the lower threshold is determined by the spatial-temporal sensing capability of the LiDAR. To assess the detection model’s performance across different datasets and parameters, we utilize high-fidelity drone and LiDAR simulations within a scenario-based Digital Twin. The selected time window affects the number of points that can be returned from an object within the LiDAR field of view. The heterogeneous LiDAR scanning pattern generates a non-uniform directivity response. This affects the object detection performance, with significant variation in the number of laser-returns backscattered by objects observed at different locations within the LiDAR field of view. The directivity response is a key factor in pre-mission risk assessment as drone observations with sparse point clouds will register as false negatives. Fig. 8 and 9 illustrate the directivity pattern of LiDAR scanning with different time windows and sensing thresholds. The field of view of the LiDAR is discretized into cubic voxels of 1.0m side length. The drone mesh model (see Fig. 4(b)) is placed at centre of each voxel. These voxels are then ray-traced over a time-windowed interval. Voxels that do not achieve the required number of ray intersections are excluded from the visualization. Only the center points of voxels that surpass the threshold are visualized and colored based on the number of intersections, i.e., generated points.
5.2 Metrics and Error Budget for Model Test
To evaluate the performance of the trained model we use the recall, precision, and F1 scores ($F1$). Recall ($R$) is the ratio of the number of true positive ($TP$) examples to the sum of true positives and false negatives ($FN$) examples. It measures the model’s ability to correctly identify all positive examples out of all the positive examples in the dataset. In the sense and detect use case it is essential to minimize the number of false negatives to ensure the safety of the system and reduce risk of accident due to collision. False negatives mean that an airborne hazard is not detected, leading to a potential mid-air collision.
@@ 5.2 Metrics and Error Budget for Model Test
Precision ($P$) is the ratio of the number of true positive ($TP$) examples to the sum of true positives and false positives ($FP$) examples. It measures the model’s ability to identify only the relevant positive examples out of all the positive predictions. In a sense and detect scenario, false positives occur when the system detects an object or obstacle that is not actually present, leading to unnecessary actions or alerts. Minimizing false positives is important because it can help reduce the number of unnecessary actions or alerts, which lead to confusion, potentially dangerous actions, and inefficiencies.
@@
We use the following thresholds for IoU measurements:
[…]
While the drones can be required to fly a path that risks denial of GNSS and communications due to multi-path and signal occlusion by large civil engineering infrastructure, a tracking-by-detection algorithm is incorporated into the labelling. The algorithm begins with an initial set of detections performed by the deep learning-based detection algorithm. These detections correspond to objects within the scene and serve as the starting point for subsequent tracking. Bounding boxes are associated with these detected objects, encapsulating each object of interest. The approximate speed of the sentry drone which is equipped with LiDAR for monitoring, is estimated by analyzing the positional change between two consecutive detections. In each successive frame, the position of tracked objects is estimated by re-centering the bounding box to the centroid of a sparse point cluster, which represents the object’s current location. If there are no points inside the re-centered bounding box, the frame is skipped, and the algorithm proceeds to the next frame. The speed information is used for the re-centering process. In practice, the deep learning-based detection algorithm and tracking algorithm are dynamically used. If deep learning-based detection successfully identifies objects in the current frame, we proceed to the next frame. Otherwise, we switch to using a tracking algorithm for that frame. The speed estimation of the sentry drone is updated using the latest two consecutive detections. The integration of tracking with detection aims to minimize both false positives (incorrectly identified objects) and false negatives (missed detections). The results of the strategy are analyzed in the results section.
6 Hardware design and system integration
6.1 System Design Requirement
For real-time sense and detect of drone in a safety scenario, the system design must adhere to the following requirements:
  • Weight and Power Limitations: The design must accommodate the constraints of drone payload capacity and battery power.
6.2 System Integration
Figure 10: Our real-time embedded system for object detection
Figure 11: Illustration of the real-time system schematic
While a higher number of point clouds can boost detection accuracy, low latency is essential for real-time hazard response. According to [51], the Avia LiDAR is capable of covering most of the field of view (FOV) in 0.1 seconds (100 milliseconds). On the other hand, our experiments demonstrate that a latency of 100ms achieves an effective balance between two competing factors: ensuring sufficient accuracy and delivering a timely response for safety. Through experimentation, we determine that the system’s maximum detection range is roughly 68m. To comprehensively assess the accuracy of our drone detection system in safety-critical scenarios, our evaluation considers three key metrics: precision, recall, and F1-score.
In safety-critical applications, the system must be independent of external aiding systems like GNSS positioning. Since GNSS multipath errors (bouncing off buildings, tree and ground etc.) may cause positioning errors such as GPS drift. By deploying the detection model directly on the drone, we eliminate the need for external aiding systems.
The designed system provides accurate 3D localization to determine the separation distance between drones within a swarm. The system successfully alerts the remote pilot if two drones come so close to each other as to violate the minimum separation rule of the operating theatre. While strategic conflict management is achieved through mission planning and pre-flight risk management, this level of tactical deconfliction is required as drones often need to deviate from pre-planned flight paths in response to an incursion by uninvolved persons, or an unscheduled return to home event due to battery depletion etc.
7 RESULTS
7.1 Performance of Sparse Convolution with Scatter Operation
We use different experiments to evaluate the proposed sparse convolution with scatter operation. An NVIDIA Jetson Xavier (32GB SSD) and JetPack 5.0.2 with CUDA 11.4 is used for the performance evaluation. The original PointPillars implementation [52] utilizes cuDNN’s dense convolution for the backbone layer. We compare the performance of our sparse convolution approach with that of the cuDNN dense convolution. As shown in Fig 12, our sparse convolution demonstrates significant performance improvements compared to the dense convolution in cuDNN. Notably, as the number of input channels increases, the performance gains become more significant. This is possibly because, when the number of input channels is small, the scattering step dominates the processing time, and its reliance on atomicAdd() operations results in slower performance. However, as the number of input channels grows, the matrix multiplication step becomes the dominant factor in processing time, benefiting from the efficient implementation provided by cuBLAS.
7.2 Model Performance Across Different Use Cases
To evaluate the sense and detect system, the embedded model was tested during supervised flights at Newcastle Aerodrome, Ireland (Fig. 13) and Hamburg Port, Germany (Fig. 14). The two test sites comprise a mix of built and natural environment and provide distinct background scenery from one another. The Newcastle site is a coastal railway bridge over a tidal inlet and is surrounded by rocks, gravel, and sand dunes. The Hamburg site is the Harburg Lock, which is a navigable waterway system that consists of a concrete chamber with gates at either end, surrounded by buildings and trees. Maintenance inspection surveys were performed at both sites involving multiple drones deployed with camera and LiDAR sensors to collect point and pixel datasets for structural analysis. The flight profiles at both sites were also significantly different, in particular the perspective and relative vantage points from which drones were observed in the LiDAR training data. This is significant as the detection model decorates input points to encode the shape, appearance, and relative position of observed objects, while the grid and pillar layout encodes the spatial relationship between the object and its surrounding environment.

### v3

@@ 4.2 Data Augmentation
4.3 Training Parameter
The parameters for the training process are as follows. For a single GPU, we limit the batch size to 4 due to memory constraints. However, when training with 8 GPUs, we increase the batch size to 32. During our experiments, we noticed that after 128 epochs, the accuracy did not improve further. Regarding the learning rate, we deviate from the standard PointPillars paper (which used 0.0001) and set it to 0.001. Our empirical observations indicate that this value is more suitable for our use case. Safety is critical in our use case scenario, so we avoid any point cloud sampling. Otherwise, small objects like drones might be excluded from the scene. In our use case, we have not encountered any situations where sampling out occurs when we set the maximum number of points each pillar to 100 and the maximum number of pillars to 12,000. However, these two values are changeable, dependent on the use case. To determine the optimal threshold values for positive IoU and negative IoU, we conducted numerous experiments with various combinations. Our observations indicate that the combination of 0.4 for positive IoU and 0.35 for negative IoU yields the highest accuracy.
5 Model Test
5.1 Model Test with Digital Twin
The LiDAR point data is accumulated into time windowed frames and then input to the detection model. The optimal time window depends on several factors. The upper threshold is determined by the acceptable computational latency which increases in proportion to the number of points in each frame, while the lower threshold is determined by the accuracy of the DNN. To assess the detection model’s performance across different datasets and parameters, we utilize high-fidelity drone and LiDAR simulations within a scenario-based Digital Twin. The selected time window affects the number of points that can be returned from an object within the LiDAR field of view. The heterogeneous LiDAR scanning pattern generates a non-uniform directivity response. This affects the object detection performance, with significant variation in the number of laser-returns backscattered by objects observed at different locations within the LiDAR field of view. The directivity response is a key factor in pre-mission risk assessment as drone observations with sparse point clouds will register as false negatives. Fig. 7 and 8 illustrate the directivity pattern of LiDAR scanning with different time windows and sensing thresholds. The field of view of the LiDAR is discretized into cubic voxels of 1.0m side length. The drone mesh model (see Fig. 3(b)) is placed at centre of each voxel. These voxels are then ray-traced over a time-windowed interval. Voxels that do not achieve the required number of ray intersections are excluded from the visualization. Only the center points of voxels that surpass the threshold are visualized and colored based on the number of intersections, i.e., generated points.
5.2 Metrics and Error Budget for Model Test
To evaluate the performance of the trained model we use the recall, precision, and F1 scores. The recall, precision, and F1 score are calculated using IoU (Intersection over Union) scores. The definitions of each metric can be found in Appendix B.
In the sense and detect use case it is essential to minimize the number of false negatives to ensure the safety of the system and reduce risk of accident due to collision. False negatives mean that an airborne hazard is not detected, leading to a potential mid-air collision. Conversely, false positives occur when the system detects an object or obstacle that is not actually present, leading to unnecessary actions or alerts. Minimizing false negatives is important because failing to detect real threats can lead to severe consequences, such as crash, rather than just minor inconveniences.
We use the following thresholds for IoU measurements:
[…]
While the drones can be required to fly a path that risks denial of GNSS and communications due to multi-path and signal occlusion by large civil engineering infrastructure, a tracking-by-detection algorithm is incorporated into the labelling. Tracking has been incorporated into point cloud-based machine learning workflows previously as in [59][60]. We utilize a simplified approach of [59], who employ a Kalman filter to predict the updated location and yaw of rigid bounding boxes for objects such as cars. In our approach, we simply re-center the bounding box from the previous frame around the enclosed points within the new frame and estimate the yaw angle. The algorithm begins with an initial set of detections performed by the deep learning-based detection algorithm. These detections correspond to objects within the scene and serve as the starting point for subsequent tracking. Bounding boxes are associated with these detected objects, encapsulating each object of interest. The approximate speed of the monitoring drone which is equipped with LiDAR, is estimated by analyzing the positional change between two consecutive detections. In each successive frame, the position of tracked objects is estimated by re-centering the bounding box to the centroid of a sparse point cluster, which represents the object’s current location. If there are no points inside the re-centered bounding box, the frame is skipped, and the algorithm proceeds to the next frame. The speed information is used for the re-centering process. In practice, the deep learning-based detection algorithm and tracking algorithm are dynamically used. If the deep learning-based detection successfully identifies objects in the current frame, we proceed to the next frame. Otherwise, we switch to using a tracking algorithm for that frame. The speed estimation of the monitoring drone is updated using the latest two consecutive detections. The integration of tracking with detection aims to minimize both false positives (incorrectly identified objects) and false negatives (missed detections). The results of the strategy are analyzed in the results section.
6 Hardware design and system integration
6.1 System Design Requirement
For real-time sense and detect of drones in a safety scenario, the system design must adhere to the following requirements:
  • Weight and Power Limitations: The design must accommodate the constraints of the drone’s payload capacity and battery power.
6.2 System Integration
Figure 9: Our real-time embedded system for object detection.
Figure 10: Illustration of the real-time system schematic.
While a denser point clouds can boost detection accuracy, low latency is essential for real-time hazard response. According to [61], the Avia LiDAR is capable of covering most of the field of view (FOV) in 0.1 seconds (100 milliseconds). On the other hand, our experiments demonstrate that a latency of 100ms achieves an effective balance between two competing factors: ensuring sufficient accuracy and delivering a timely response for safety. Through experimentation, we determine that the system’s maximum detection range is roughly 68m. To comprehensively assess the accuracy of our drone detection system in safety-critical scenarios, our evaluation considers three key metrics: precision, recall, and F1-score.
7 Performance of Sparse Convolution with Scatter Operation
TABLE II: Comparison between our model and PCDet models.
3D Detection Model Modification Precision (%) Inference time (ms)
Centerpoint-PillarNet Extend Z-axis range 89.38 34.13
DSVT-PillarNet Extend Z-axis range 67.86 62.25
Our model (modified PointPillars) Extend Z-axis range only 83.53 14.16
Extend Z-axis range + our (sparse+submanifold) convolution 82.60 6.13
We use different experiments to evaluate the proposed sparse convolution with scatter operation. An NVIDIA Jetson Xavier (32GB SSD) and JetPack 5.0.2 with CUDA 11.4 is used for the performance evaluation. The standard PointPillars implementation [62] utilizes cuDNN’s dense convolution for the backbone layer. We compare the performance of our sparse convolution approach with that of the cuDNN dense convolution. As shown in Fig 11, our sparse convolution demonstrates significant performance improvements compared to the dense convolution in cuDNN. Notably, as the number of input channels increases, the performance gains become more significant. This is possibly because, when the number of input channels is small, the scattering step dominates the processing time, and its reliance on atomicAdd() operations results in slower performance. However, as the number of input channels grows, the matrix multiplication step becomes the dominant factor in processing time, benefiting from the efficient implementation provided by cuBLAS.
8 Model Performance Across Different Use Cases
8.1 Experiment 1: Newcastle use case
Figure 12: Schematic diagram for dataset generation and model training.
TABLE III: Evaluation of the detection model on Newcastle dataset
model_sim (data_sim)(data\_sim) only applied 0.511 0.737 0.603
not applied 0.949 0.788 0.861
(data_sim)(data\_sim) + (real_first_400_frame)(real\_first\_400\_frame) applied 0.523 0.75 0.616
not applied 0.96 0.80 0.872
model_euc (data_euc)(data\_euc) + (real_first_400_frame)(real\_first\_400\_frame) applied 0.24 0.774 0.367
not applied 0.766 0.513 0.614
Figure 13: Visualization of predicted bounding box (red) and ground truth bounding box (green), the inference model used is (model_sim)(model\_sim) described in Section 8.1, a portion of the frame containing the drone is enlarged for clarity. (a) Newcastle scenario: Two drones deployed to survey the railway bridge. The red arrow indicates the target drone in the image captured by the camera of the monitoring drone. The small picture inside the image is captured by the high resolution camera of photogrammetry drone. (b) The predicted bounding box has a good overlap with the ground truth bounding box (c) The predicted heading is different with the ground truth (d) The difference between the center of predicted bounding box and the center of ground truth bbox is significant.
The objective in this experimentation with the Newcastle scenario was to evaluate the performance of the model and the data augmentation strategy. The data generation scheme and model training are depicted in Fig. 12, with a detailed explanation provided in Appendix C. The training, validation, and testing split ratio is 2.4:1:2.2.
8.1.1 Performance on real data
We first train our 3D object detection model on real data without augmentation and compare its performance to detection models available in PCDet [63], an open-source toolbox designed for point cloud-based 3D object detection. PCDet is widely used for LiDAR-based perception tasks and supports multiple state-of-the-art detection methods. Its codebase is highly modular and includes both voxel-based and pillar-based models. For our comparison, we focus on pillar-based models from PCDet, as they do not impose strict restrictions on the Z-axis range, unlike voxel-based models. This flexibility allows us to adapt these models for drone-captured data with varying altitudes. Specifically, we compare our model with the state-of-the-art CenterPoint-PillarNet and DSVT-PillarNet models from PCDet, modifying their Z-axis range to better accommodate our drone data with varying altitudes. CenterPoint model reformulates detection as a center-based approach rather than relying on anchors or dense bounding box predictions. DSVT is an attention-based model that selectively processes only the important sparse voxel regions. Table II presents a comparison between our model and the PCDet models, using a single NVIDIA RTX 3090 GPU. We used 1,603 real frames from the Newcastle dataset for training.
As shown in Table II, CenterPoint-PillarNet achieves the highest precision but at the cost of slower inference speed compared to our model. DSVT-PillarNet performs poorly in both precision and inference time. Our model delivers the fastest inference speed. Specifically, the version of our model with both Z-axis range extension and (sparse+submanifold) convolution acceleration achieves the fastest inference time, with only a slight precision sacrifice (less than 1%) compared to the model with Z-axis range extension alone. In our safety-critical application scenario, response time is crucial for real-time hazard avoidance. Therefore, we selected and modified PointPillars to prioritize faster inference times while maintaining acceptable precision.
8.1.2 Inference time on embedded hardware
We evaluated the inference runtime on the Newcastle dataset. As shown in Table IV, for inference with FP32, the Pillarization part (Pillar Feature Net) accounts for approximately 1.8% of the total inference time. The Backbone + Head part is the most time-consuming, making up 81.8% of the overall runtime. The Postprocess + NMS (Non-Maximum Suppression [64]) step contributes around 16.4% to the total inference time.
Regarding precision, the mixed precision implementation significantly reduced the overall runtime, particularly in the Backbone + Head. This shows that optimization through mixed precision making it a promising approach for real-time applications on embedded hardware like the Jetson Xavier.
TABLE IV: Runtime of each model component on NVIDIA Jetson AGX Xavier.
Model Component
Runtime (ms)
Runtime (ms)
(Mixed Precision)
Pillarization 1.21 0.82
Backbone + Head 54.27 32.49
Postprocess + NMS 10.89 8.16
Overall 66.37 41.47
8.1.3 Performance on augmented data
As shown in Table III, the model trained on a point cloud consisting of both synthetic augmentation data (data_sim)(data\_sim) and real data from the first 400 frames (real_first_400_frame)(real\_first\_400\_frame), without utilizing intensity values, achieves the best performance in all of the evaluation metrics, Precision, Recall, and F1 score. However, the model’s performance decreases when intensity values are used for training, indicating that there is a discrepancy between the the simulated intensity values and the ground truth intensity values. The model trained on a point cloud generated by combining Euclidean augmentation of real data, specifically (data_euc)(data\_euc) and (real_first_400_frame)(real\_first\_400\_frame), without utilizing intensity values, achieves a competitive performance. The model’s performance decreases when using intensity values for training, potentially due to the aforementioned possible discrepancy. The model trained solely on simulated data (data_sim)(data\_sim) shows a slight decrease in all metrics compared to the model trained on a combination of (data_sim)(data\_sim) and (real_first_400_frame)(real\_first\_400\_frame). This demonstrates that the model trained with simulated data alone can achieve performance remarkably close to that of the model trained with both simulated and real data. The model trained on real data without any modifications achieves a comparable performance. However, it produces more false negatives, suggesting a lower capability for detecting drones. These results demonstrate that our simulation augmentation strategy can generate more accurate data than Euclidean augmentation of real data.
Visualization of both the predicted bounding box and the ground truth bounding box is shown in Fig. 13, where the inference model applied is (model_sim)(model\_sim). Through our testing, we discovered that predicting the heading of the detected drone is the most challenging task. This could be attributed to the fact that heading prediction typically requires a sufficient number of points representing the geometric structure of the body of the drone. However, in most cases, the LiDAR cannot generate an adequate number of points, because the drone is such a small object. In certain instances, as depicted in Fig. 13(d), there is a noticeable difference between the center of the predicted bounding box and that of the ground truth bounding box. This observation suggests that our model is imperfect.
8.2 Experiment 2: Hamburg use case
This test involves an inspection survey with two drones (see the top left part of Fig. 14) simultaneously flying LiDAR (circled red) and camera (circled green) data capture missions. The detection model was trained on data that was previously captured in the lock and augmented with synthetic observations using the digital twinning method described earlier. The 100ms point cloud scenes presented to the detection model embedded on the monitoring drone typically contained observations of the LiDAR drone positioned relatively high and well clear of other objects, as well as the camera drone positioned relatively low and near the concrete wall.
The sense and detect system combined the deep learning based detection model with the tracking-by-detection algorithm described earlier. The lower part of Fig. 14 displays a graph of the distance between the two drones during an example close proximity fly-by. The quantitative data is presented in Table V. There are 430 frames in total, representing 43 seconds of flight perception data and translating to an average relative speed of c. 1m/s between the two drones.
Figure 14: Detection results at Harburg Lock. Top left: Harburg Lock scenario: Two drones (circled red and green) deployed for an inspection survey, the third drone is (circled yellow) monitoring their separation distance. Top right: Detection of two drones using the sense and detect system. The red bounding box is predicted by the deep-learning based detection model and the green bounding box is by the tracking algorithm. Bottom: Distance between the two drones fly-by, where the distance is computed using the detection results of detection and tracking algorithm.
TABLE V: Quantitative data associated with graph in Fig. 14.
Two drones tracked by inference model detection 333
Two drones tracked by combination of inference model (1)
and tracker algorithm (1)
9
Only one drone tracked by either method 83
No drones detected or tracked 5
9 Discussion
9.1 Detection Performance and Real-World Feasibility
This work presents the first airborne LiDAR-based deep learning detection framework designed to operate entirely onboard an airborne embedded platform, enabling real-time detection without reliance on ground-based processing. The results demonstrate that our model achieves high recall and precision in airborne conditions, even in complex operational environments typical of transport system infrastructure. Our findings show that incorporating Digital Twin-based augmentation significantly improves detection robustness, with models trained on augmented data achieving a recall improvement of 18% and precision improvement of 13% compared to conventional augmentation techniques. The ability to enhance model generalization through physics-informed augmentation is crucial for real-world deployment, where training data diversity is often limited.
9.2 Sparse Convolution Acceleration and Embedded Efficiency
A major challenge in deploying deep learning-based airborne detection is the severe computational constraints imposed by embedded systems. Unlike automotive applications, where dedicated high-power GPUs are available, airborne platforms must balance detection accuracy with real-time inference speed while operating under strict size, weight, and power (SWaP) constraints. Our scatter-based sparse convolution approach addresses this challenge, reducing inference latency by 2.3× while maintaining competitive accuracy. The observed efficiency gains scale with increasing input channels, as the optimized implementation shifts the computational burden toward matrix multiplication, benefiting from cuBLAS acceleration. This result confirms that real-time LiDAR-based perception is feasible onboard airborne systems without requiring specialized hardware accelerators.
9.3 Challenges in Deploying LiDAR Detection in Constrained Environments
The field deployments highlighted several practical challenges for airborne LiDAR-based detection in real-world infrastructure monitoring. The detection performance was affected by targets frequently appearing at the periphery of the LiDAR scan region, reducing effective point density. Structural elements introduced multipath and erroneous returns, requiring additional filtering and tracking techniques. Unlike static infrastructure, the presence of moving vessels, water reflections, and confined spaces challenged the detection consistency. These results indicate that while airborne LiDAR detection is feasible, some adaptations are required for deployment in enclosed or highly dynamic operational settings.
10 Conclusion
This paper tackles the challenge of monitoring a drone swarm during collaborative tasks. It presents the first airborne LiDAR based solution for drone-swarm detection and localization using a 3D deep learning model. It adapts an existing detection model, originally designed for ground applications, to an air-to-air drone scenario by expanding the scan space vertically. To reduce the latency of the detection model, we propose a new sparse convolution algorithm that improves efficiency by bypassing the mapping between input and output, which is required by existing methods. This algorithm accelerates the model’s backbone layer, which is the most time-consuming component of the detection process. A scenario Digital Twin is implemented to simulate close encounters and safety-critical scenarios to gather training data. By augmenting the actual background data with high-quality synthetic drone data, we improve the accuracy and efficiency of our training and inferencing processes. Moreover, we incorporate a tracking-by-detection algorithm into the system, which allows for accurate monitoring of the separation distance of multiple drones, even in challenging environments.
Our approach has been successfully validated through real-world tests, with the trained model achieving an impressive 80% recall and 96% precision when tested on real datasets. We trained the detection model on several datasets, each generated using different augmentation strategies, and found that our synthetic augmentation strategy produced more precise data than the conventional augmentation method used with real data. Overall, our proposed approach offers a novel and effective solution for drone swarm detection and localization, with promising results from our tests.
Appendix A LiDAR simulation with NVIDIA Optix Ray Tracing
@@
  • Step 2. Move the origin of the coordinate system to the center of the drone mesh model.
@@
Each OptiX thread computes a new ray origin and direction vector independently, rather than rebuilding the acceleration structure in the ray tracing core.
@@
Appendix B Evaluation Metrics
Recall ($R$) is the ratio of the number of true positive ($TP$) examples to the sum of true positives and false negatives ($FN$) examples. It measures the model’s ability to correctly identify all positive examples out of all the positive examples in the dataset.
@@ Appendix B Evaluation Metrics
Precision ($P$) is the ratio of the number of true positive ($TP$) examples to the sum of true positives and false positives ($FP$) examples. It measures the model’s ability to identify only the relevant positive examples out of all the positive predictions.
@@
Appendix C Data generation scheme
We collect training data by scanning the target drone (detection object) using a LiDAR sensor mounted on a monitoring drone. This process begins with the target drone positioned 10 meters away from the monitoring drone, after which it flies along the x-axis. It is important to note that the LiDAR sensor serves as the origin of the 3D coordinate system and is oriented along the x-axis. As demonstrated in the paper, the non-uniform scan pattern of the LiDAR sensor results in the highest point density when the target drone moves along the x-axis. Once the target drone reaches a given point, it will change elevation.
The collected LAS dataset (data_orig)(data\_orig) was segmented into 100ms frames, and the target drone observations were labelled with bounding boxes placed at locations derived from the target drone flight log. The dataset was then reduced into the set of frames (N=1603) that contained bounding boxes enclosing at least 10 points. The first 400 instances (real_first_400_frame)(real\_first\_400\_frame) of the resulting frames were filtered to extract observations of the target drone and the background separately. We call them (extracted_drone)(extracted\_drone) and (extracted_background)(extracted\_background) respectively. During data collection, these 400 frames were recorded while the target drone moves along the x-axis. This approach ensures that each frame contains the densest point clouds, effectively preserving the geometric shape of the drone from various positions. On the other hand, the first 400 frames are recorded right after IMU calibration. Frames recorded right after IMU calibration will have the tightest positional accuracy for labelling. Frames could be chosen from later in the dataset after later IMU calibrations, but the first 400 are the easiest timestamp wise to utilise. By augmenting these 400 frames, we can generate a high-quality, geometry-preserving augmented dataset, improving the robustness and reliability of the training data.
Our augmentation strategy simulates LiDAR scans by placing a 3D drone mesh at randomized orientations within a stratified FOV grid and performing ray tracing over 100ms to generate physically accurate point clusters, retaining only those with at least ten valid hits. Each of the 2495 instances was then randomly matched to one of the 400 backgrounds (extracted_background)(extracted\_background). The resulting dataset (data_sim)(data\_sim) and the real 400 frames (real_first_400_frame)(real\_first\_400\_frame) were used together to train the detection model (model_sim)(model\_sim). The set of insertions locations was retained and reused in generating a comparative dataset from Euclidean augmentations in the second strategy.
To generate the Euclidean augmentation dataset (data_euc)(data\_euc), each instance of (extracted_drone)(extracted\_drone) were upsampled into 2895 exemplars by creating 2495 additional rotated copies. Each point cluster was then translated to the corresponding background and insertion locations retained during construction of the data_simdata\_sim dataset. The resulting dataset (data_euc)(data\_euc) was used to train the detection model (model_euc)(model\_euc).
Both training datasets are the same size, contain the same number of augmentations, and corresponding frames have the same background with a drone inserted in the same location. To test the performance of the respective models, we collected new LiDAR data from the same mission site, which consists of consecutive 2694 frames. In these 2694 frames, drones (ground truth) observed in 1555 frames.
Appendix D Survey sites
The data specification of maintenance inspection surveys facilitates swarm flights as multiple drones can simultaneously fly spatially segmented waypoint- and route-based data collection missions as the different point- and pixel- datasets have different resolution requirements. For example, collection of sub millimeter Ground Sample Density camera imagery in photogrammetry and crack detection use cases can require the drone to fly within 2-6m of the target structure etc., while LiDAR surveys aiming for centimeter point density can fly at offset distances of 20-30m. This facilitates separation and deconfliction of planned swarm trajectories during strategic flight risk management.
D.1 Newcastle scenario
The Newcastle scenario see Fig. 15 involved two drones deployed to survey the railway bridge, one drone carrying a LiDAR sensor and the other with a high-resolution camera for photogrammetry. The LiDAR flight was approximately 10 minutes in duration, including IMU calibrations at the start and end of the mission. The LiDAR drone frequently observed the camera drone, with observations in the point cloud data annotated using positional information extracted from the flight logs.